OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4)

OpenStudy (nurali):

slope-intercept form; y=mx+b Lets start by using point-slope form; (Y-Y1)=m(X-X1) (x1,y1)=(3, -4) and slope=-3 Now plug in the information given; (y+4)=-3(x-3) y+4=-3x+9 y=-3x+9-4 y=-3x+5

OpenStudy (nurali):

Welcome to Openstudy @riahhroo

OpenStudy (anonymous):

thank you!

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (nurali):

i will try.

OpenStudy (anonymous):

Write the equation of a line in point slope form that has a slope of 4 3 − and passes through the point (4, 5).

OpenStudy (anonymous):

-3 over 4 *

OpenStudy (nurali):

slope-intercept form; y=mx+b Lets start by using point-slope form; (Y-Y1)=m(X-X1) (x1,y1)=(4, 5) and slope=-3/4 Now plug in the information given; (y-5)=-3/4(x-4) 4y-20=-3x+12 4y=-3x+12+20 4y=-3x+32